What is the moment of inertia of a 10 kg stone traveling along a straight line?
Answer #1

Let the stone shape is solid spherical of radius r

Hence moment of inertia = I = 2/5 mr^2 = = 2/5 × 10r^2 = 20r^2

=) I = 20r^2 kg.m^2

Here you didn't mention radius. Please put the value to radius r t get numerical value.